OpenStudy (goformit100):

The number of values of k, for which the system of equations (k + 1)x + 8y = 4k kx + (k + 3)y = 3k − 1 has no solution, is (1) 1 (2) 2 (3) 3 (4) infinite

OpenStudy (nurali):

i think 2

OpenStudy (anonymous):

For them to have no solution, the lines described by the equations must be parallel, i.e. $$k+1=kt\\8=(k+3)t\\4k=3kt$$How many \(k\) satisfy the above constraints for arbitrary \(t\)?

OpenStudy (anonymous):

$$4k=3(k+1)\\4k=3k+3\\k=3\\t=\frac43$$... this is the only solution I can find.

OpenStudy (anonymous):

It's an overdetermined system anyways.